Update - We are aware of the remaining issues. Our engineers are doing everything they can to return service to all customers. Once everything is resolved we will update you here.
Mar 17, 11:36 UTC
Update - Service is starting to restore. You should be able to access servers at this point. We are still monitoring the system, and performance will be degraded for the time being.
Mar 17, 10:09 UTC
Update - Our engineers are still on site working hard to resolve this issue, we will not sleep until this has been resolved.
Mar 17, 08:15 UTC
Monitoring - We are still working on this issue, and there is still no ETA on when this will be resolved. We have a team of engineers on site already looking into resolving the issue. Again, we apologize to our affected customers, but we will provide another update as soon as we have one.
Mar 17, 02:31 UTC
Update - We are working directly with our vendor to resolve the issue. At this time we think this could be a possible bug in their system and something only they can fix. We will post updates here as soon as we have more information.
Mar 17, 00:23 UTC
Update - Our engineers are still looking into this. We will provide an update in 1 to 2 hours.
Mar 16, 21:23 UTC
Investigating - We are investigating a SAN issue with our cluster01 environment. We have all engineers working on this matter. We thank you for your patience.
Mar 16, 19:06 UTC